OpenStudy (anonymous):

220g of Co2 are heated from 27 C - 527C at a constant pressure 1 atm and then compressed isothermally to 100atm calculate the energy absorbed in K.cal Assume Ideal Behaviour (cp=5 cal deg^-1 mol^-1)

OpenStudy (anonymous):

@Vincent-Lyon.Fr @akash123 @Carl_Pham

OpenStudy (anonymous):

\[\Delta U = W + Q\] We have to Find Q right ?

OpenStudy (anonymous):

Or Is there any other Way to Appraoch This,,,,,)

OpenStudy (anonymous):

yes...u have to do this only...calculate work done and change in internal energy

OpenStudy (anonymous):

\[\Delta U = nCv \Delta T\]

OpenStudy (anonymous):

But i am Confussed Hw to Find Work..)

OpenStudy (anonymous):

work done in isothermal process w= 2.303 RT log(v2/v1)

OpenStudy (anonymous):

use w= integral of Pdv substitute pp from pv=nRT, u'll get the formula for work done in isothermal process.

OpenStudy (anonymous):

But V2 and V1.... ???

OpenStudy (anonymous):

use pv=n RT

OpenStudy (anonymous):

w= 2.303 RT log(p1/p2).???

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

P1 = 1 P2 = 100

OpenStudy (anonymous):

T = 527 C

OpenStudy (anonymous):

wait...there r two parts in this Q

OpenStudy (anonymous):

in 1st part...it's isobaric process so simply work done=p(v2-v1)

OpenStudy (anonymous):

2nd process is isothermal one

OpenStudy (anonymous):

so total work done= w(isobaric)+ w(isothermal)

OpenStudy (anonymous):

Ok....For Isothermal.... T= 527....and P1 = 1 P2 =100

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

use pv=nRT for finding v1 and v2 in the isobaric process.

OpenStudy (anonymous):

For..Lemme Try...and will Inform u...) thxxx....

OpenStudy (anonymous):

fine...:)

OpenStudy (anonymous):

Lol....i got 15.4 K. cal...But the answer shuld be 19.35 K. cl @akash123

OpenStudy (anonymous):

Can u check..) pLzzz

OpenStudy (anonymous):

U hv calculated total work done only...but u have to calculate change in internal energy also.

OpenStudy (anonymous):

du = n c(v) dT

OpenStudy (anonymous):

I have Done...that too...

OpenStudy (anonymous):

Is Cv = 3

OpenStudy (anonymous):

what did u use for c(v)? c(v) =1/2 f R f is the degree of freedom so CO2 is a triatomic so how many degree of freedom for CO2?

OpenStudy (anonymous):

Cp - Cv = R Cp - R = Cv

OpenStudy (anonymous):

R=~2

OpenStudy (anonymous):

but how did u come to know the value of c(v) without knowing f?

OpenStudy (anonymous):

Cp is Given..)

OpenStudy (anonymous):

Meyar"s Relation cp-cv=R

OpenStudy (anonymous):

but u know c(p)?

OpenStudy (anonymous):

oh...sorry...den u r right.

OpenStudy (anonymous):

then for isothermal...change in internal energy=0 u have to calculate only for isobaric process....how much u got?

OpenStudy (anonymous):

5000 Cal.

OpenStudy (anonymous):

P V1 = nRT V1 = 5 * 2 * 300

OpenStudy (anonymous):

and w(isobaric) and w(isothermal)?

OpenStudy (anonymous):

look for the units carefully...

OpenStudy (anonymous):

for ex V1 = 5 * 2 * 300 p--atm R---cal/mol/K

OpenStudy (anonymous):

so u'll get V in cal/atm

OpenStudy (anonymous):

For an ideal gas U depends only on T via the energy equation of state, U = C_v NRT, and you can use C_V = C_p - R here since you're given C_p. I don't see a need to calculate the work and heat.